OpenStudy (anonymous):

A rich aunt tells you that you can choose one of the following: A. During a period of 15 days, she will give you 1 cent the first day, give you double that the second day, give you double the second on the third day and so on. B. She will give you $100. Which option would give you more money, A or B?

OpenStudy (anonymous):

As a general rule of thumb, the answer is ALWAYS the one that doubles :-P

OpenStudy (anonymous):

A. 1 +2^1 + 2^2 + 2^3 +........+ 2^14 = S 1 + 2 x (1 +2^1 + 2^2 + 2^3 +........+ 2^13) = S 1+ 2 x ( S - 2^14) = S 1+ 2S - 2^15 = S S = 2^15 - 1= 32767 cents = 327.67 $ B. 100 $

OpenStudy (anonymous):

A.